OpenStudy (anonymous):

Let f(x) = x2 + 6 and g(x) = x + 8 divided by x. Find (g ° f)(–7).

OpenStudy (anonymous):

\[ (f\circ g)(-7) = f(g(-7)) = f((-7)+8) = [(-7)+8]^2+6 \]

OpenStudy (anonymous):

@mathhelp7 Does this make sense to you?

OpenStudy (anonymous):

No, how would I find the answer?